Question 4

Let f be the function defined by the formula,

f(x) = 1/x+1/x − 10

.

a) Determine the largest possible domain D of f.

b) Is f injective on D?

[8,5]


Question 5

Compute the f ◦ g and its range of the functions f and g below,

f(x) = (x^2 + 5x − 6)(x^2 + 5)/|2x + 3|

, and g(x) = √x + 4

[12]


Question 6

Determine the largest domain, intersection with axes, and sign of f

f(x) = log2(2 −2/x − 3)

[16]


1
Expert's answer
2021-10-18T16:07:32-0400

4.

a)

"D=\{x\isin (-\infin,0)\lor (0,10)\lor (10,\infin)\}"


b)


an injective function is a function f that maps distinct elements to distinct elements; that is,

"f(x_1)=f(x_2)" implies "x_1=x_2"

So, f(x) is injective.


5.

"f\circ g=\frac{(6x-2)(x+9)}{|2\sqrt{x+4}+3|}"


Range of f(x): "(-\infin,\infin)"

Range of g(x): "[0,\infin)"

for "f\circ g" :

"x\isin [-4,\infin)"

"f\circ g(-4)=\frac{-22\cdot 5}{3}=-\frac{110}{3}"

Range of "f\circ g" : "[-110/3,\infin)"


6.

for f(x):

"2-\frac{2}{x-3}>0\implies \frac{2x-8}{x-3}>0"

domain: "x\isin (-\infin,3)\lor (4,\infin)"

for x-intersection:

"2-\frac{2}{x-3}=1\implies x=5"

x-intersection: "(5,0)"

for y-intersection:

"log_2(2+2\/3)=log_2(8\/3)=3-log_23"

y-intersection: "(0,3-log_23)"

for sign of f:

"f(x)<0\implies 0<\frac{2x-8}{x-3}<1\implies |x-3|>|2x-8|"

"f(x)<0" at "x\isin (4,5)"

"f(x)>0" at "x\isin (-\infin,3)\lor (5,\infin)"